OpenStudy (anonymous):

use the communative property to identify witch expression is equal to (2=5)*8 A. (5+2)*8 B. 8*2+5*8 C. 2+(5*8) D. (2+5*8)

OpenStudy (anonymous):

help with this please? will give medal

OpenStudy (anonymous):

ITS NOT SOPOST TO BE AN EQUAL SIGN ITS SOPOST TO BE A PLUS SIGN

OpenStudy (anonymous):

luk here u need to follow d RULE of BODMAS ... i.e., Bracket,Of,Division,Multiplication,Addition,Substraction. in sequence .. so here, (2+5)*8 = 7*8=56 or, 2*8 + 5*8 = 56 (commutative)

OpenStudy (anonymous):

ohhh alright so what would the answer bee?

OpenStudy (anonymous):

tanjeetsarkar96

OpenStudy (anonymous):

Ans is B .